40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15570     Entry time: Tue Sep 15 10:57:30 2020
Author: gautam 
Type: Update 
Category: PSL 
Subject: PMC re-locked, RGA re-enabled 

The PMC has been unlocked since September 11 sometime (summary pages are flaky again). I re-locked it just now. I didn't mess with the HEPA settings for now as I'm not using the IFO at the moment, so everything should be running in the configuration reported here. The particulate count numbers (both 0.3um and 0.5um) reported is ~x5-8 of what was reported on Thursday, September 10, after the HEPA filters were turned on. We don't have logging of this information in any automated way so it's hard to correlate things with the conditions in Pasadena. We also don't have a working gauge of the pressure of the vacuum envelope.

The RGA scanning was NOT enabled for whatever reason after the vacuum work. I re-enabled it, and opened VM1 to expose the RGA to the main volume. The unit may still be warming up but this initial scan doesn't look completely crazy compared to the reference trace which is supposedly from a normal time based on my elog scanning (the timestamp is inherited from the c0rga machine whose clock is a bit off).


Update 1500: I checked the particle count on the PSL table and it barely registers on the unit (between 0-20 between multiple trials) so I don't know if we need a better particle coutner or if there is negligible danger of damage to optics due to particulate matter.

Attachment 1: RGAscan.pdf  173 kB  Uploaded Tue Sep 15 13:06:01 2020  | Hide | Hide all
RGAscan.pdf
ELOG V3.1.3-